OpenStudy (anonymous):

Factor by grouping 4r^2+9r-28

OpenStudy (anonymous):

4r2+16r-7r-28=0 4r(r+4)-7(r+4)=0 (4r-7)(r+4)=0 r = -4 or 7/4